OpenStudy (anonymous):

Show that 3x^2 - 4x + 2 is always positive, for all real values of x.

OpenStudy (zarkon):

complete the square

OpenStudy (anonymous):

can someone help me with this, Michael rolled two number cubes, numbered 1 to 6, simultaneously. According to theoretical probability, how many of the rolls should Michael expect the sum to be 6 if he rolled them 108 times?

sam (.sam.):

\[\fbox{$3 \left(x-\frac{2}{3}\right)^2+\frac{2}{3}$}\]

sam (.sam.):

Vertex at (2/3,2/3)

OpenStudy (anonymous):

Thank you again, Sam! (: You're very helpful. And a genius. Hahahaha. And thanks, Zarkon! (: